# Large Stepsize Gradient Descent for Logistic Loss: Non-Monotonicity of the Loss Improves Optimization Efficiency

### v1

@@ Abstract
We consider gradient descent (GD) with a constant stepsize applied to logistic regression with linearly separable data, where the constant stepsize $\eta$ is so large that the loss initially oscillates. We show that GD exits this initial oscillatory phase rapidly — in $\mathcal{O}(\eta)$ steps — and subsequently achieves an $\tilde{\mathcal{O}}(1/(\eta t))$ convergence rate after $t$ additional steps. Our results imply that, given a budget of $T$ steps, GD can achieve an accelerated loss of $\tilde{\mathcal{O}}(1/T^{2})$ with an aggressive stepsize $\eta:=\Theta(T)$, without any use of momentum or variable stepsize schedulers. Our proof technique is versatile and also handles general classification loss functions (where exponential tails are needed for the $\tilde{\mathcal{O}}(1/T^{2})$ acceleration), nonlinear predictors in the neural tangent kernel regime, and online stochastic gradient descent (SGD) with a large stepsize, under suitable separability conditions.
@@ 1 Introduction
The classical theory for constant-stepsize GD assumes a small stepsize such that the induced loss decreases monotonically (Nesterov, 2018). For example, if $L(\mathbf{w})$ is $\beta$-smooth, then choosing $\eta<2/\beta$ guarantees $L(\mathbf{w}_{t})$ to decrease monotonically, a result known as the descent lemma (Nesterov, 2018, Section 1.2.3). Fruitful convergence results have been developed for small stepsize GD based on the descent lemma or its variants. However, little theory is known when GD is run with a large stepsize (for instance, $\eta>2/\beta$ for a $\beta$-smooth function), which induces non-monotonic loss values (exceptions will be discussed later in Section 1.1). Cohen et al. (2020) refer to this as the edge of stability (EoS) regime, and they further point out that GD usually operates in the EoS regime for training neural networks to exhibit reasonable optimization or generalization performance.
@@
Our first main result (Theorem 1) characterizes the dynamics of large stepsize GD for logistic regression with linearly separable data. We show that GD initially induces a non-monotonic loss (the EoS phase) and then exits this regime in finite time (phase transition), and afterwards the loss decreases monotonically (the stable phase). Specifically:
@@
    The Stable phase.  Finally, we show in the stable phase, the loss decreases monotonically at an $\tilde{\mathcal{O}}\big{(}1/(\eta t)\big{)}$ rate after $t$ steps. Figures 1(b) and 1(d) empirically verify that this rate is sharp asymptotically (as $t\to\infty$) ignoring logarithmic factors.
The above result immediately justifies the benefits of a large stepsize for improving optimization efficiency. First, a larger $\eta$ yields a smaller constant factor in the asymptotic bound, $\tilde{\mathcal{O}}(1/(\eta t))$. We highlight that $\eta$ can be an arbitrarily large constant, beyond the classical wisdom that constrains $\eta$ by the inverse smoothness (Nesterov, 2018). Second, given a budget of $T$ steps, GD attains an $\tilde{\mathcal{O}}(1/T^{2})$ loss when equipped with an aggressive stepsize $\eta:=\Theta(T)$ which optimally balances the steps spent in the EoS and stable phases (see Corollary 2). In contrast, we also provide an $\Omega(1/T)$ loss lower bound for constant-stepsize GD that does not enter the EoS phase (see Theorem 3). Our theory explains the empirical observations in Figure 1.
@@
  • General losses.  We relax the logistic loss in (1) to general classification loss functions under a set of conditions (see Assumption 3 and Theorem 6). Specifically, the EoS bound relies on the Lipschitzness of the loss, the stable phase bound relies on an additional self-boundedness condition, and the $\tilde{\mathcal{O}}(1/T^{2})$ acceleration relies on the loss having an exponential tail.
  • Nonlinear models.  We extend the linear predictor in (1) to a two-layer network in the neural tangent kernel (NTK, Jacot et al., 2018) regime (Theorem 6). To our knowledge, this is the first NTK result that accommodates an arbitrarily large stepsize (although the minimum network width depends on the stepsize). For a constant stepsize, the minimum network width is polylogarithmic in the number of steps and the number of samples, recovering the main result of Ji and Telgarsky (2019) as a special case.
  • SGD.  Finally, we consider constant-stepsize online stochastic gradient descent (SGD) for logistic regression with separable data. We provide upper bounds on the population logistic loss and the population zero-one error (Theorem 4), allowing the stepsize to be $o(T)$ (where $T$ is the maximum number of steps) and arbitrarily large, respectively. These are in contrast to the typical SGD results in online smooth convex optimization that only allow small (even vanishing) stepsizes (see, for example, Hazan, 2022, Chapter 3).
(a) Training loss, toy dataset.
(b) Asymptotic rate, toy dataset.
(c) Training loss, subet of MNIST.
(d) Asymptotic rate, subset of MNIST.
Figure 1: Behaviors of (GD) for logistic regression (1) with initialization $\mathbf{w}_{0}:=0$. Figures 1(a) and 1(b) are results on a toy dataset consisting of four samples, $\mathbf{x}_{1}=(1,0.2),\ y_{1}=1$, $\mathbf{x}_{2}=(-2,0.2),\ y_{2}=1$, $\mathbf{x}_{3}=(-1,-0.2),\ y_{3}=-1$, and $\mathbf{x}_{4}=(2,-0.2),\ y_{4}=-1$. Here, the stepsize $\eta=2^{2}$ is small, and $\eta\in\{2^{3},2^{4},2^{5}\}$ is large (depending on whether the loss decreases monotonically). Figures 1(c) and 1(d) are results on $1000$ samples from the MNIST dataset with labels “0” or “8”. Here, the stepsize $\eta=2^{-6}$ is small, and $\eta\in\{2^{-4},2^{-2},2^{0}\}$ is large. Figures 1(a) and 1(c) suggest that (GD) with a larger stepsize incurs a smaller loss when stabilized. Figures 1(b) and 1(d) suggest that $L(\mathbf{w}_{t})=\Theta(1/(\eta t))$ asymptotically as $t\to\infty$.
@@ 1.1 Related Works
Empirically, training neural networks with GD (or large batch SGD) often requires a large stepsize to achieve reasonable optimization and generalization performance (see, for example, Hoffer et al., 2017; Goyal et al., 2017; Wu and Ma, 2018; Cohen et al., 2020), whereby the training loss exhibits non-monotonic behavior. Cohen et al. (2020) introduce the term edge of stability (EoS) to refer to this. The theory of EoS and large stepsize has been studied in various cases, such as one- or two-dimensional (nearly analytic) functions (Zhu et al., 2022; Ahn et al., 2023; Kreisler et al., 2023; Chen et al., 2023; Wang et al., 2023), scale-invariant networks (Lyu et al., 2022), diagonal linear networks (Even et al., 2023; Andriushchenko et al., 2023), and matrix factorization (Wang et al., 2022b; Chen and Bruna, 2023). Besides, there are some general theories for EoS (Kong and Tao, 2020; Ahn et al., 2022; Damian et al., 2022; Ma et al., 2022; Wang et al., 2022c), but they are built up on subtle assumptions. In comparison, we work in a natural classification setting under standard assumptions.
Among the literature of EoS, the work by Wu et al. (2023) is most relevant to us. They show the convergence of GD with an arbitrarily large stepsize for logistic regression with linearly separable and non-degenerate data. Our work significantly improves theirs. First, the constant factors in their bounds are exponential in the stepsize $\eta$. In comparison, we nail down the dependence on $\eta$ in each phase (see Theorem 1), establishing the $\tilde{\mathcal{O}}(1/(\eta t))$ asymptotic rate and the $\tilde{\mathcal{O}}(1/T^{2})$ acceleration with stepsize $\eta=\Theta(T)$ (see Corollary 2). So we explain the benefits of large stepsize while they cannot. Second, they require the dataset to be non-degenerate such that the support vectors span the dataset (see their Assumption 3), while we do not need this condition. Finally, their results are specialized to GD and logistic regression, while our approach is versatile and also handles general classification losses, nonlinear predictors in the NTK regime, and SGD with a large stepsize.
@@
The neural tangent kernel (NTK) is most prominently presented and named by Jacot et al. (2018), and similar concepts are used by subsequent works to analyze the early training dynamics of GD. In the majority of these works (see Du et al., 2018; Zou et al., 2018; Zou and Gu, 2019; Allen-Zhu et al., 2019, for example), the stepsize of GD is $o(1)$ and vanishes as a function of other parameters such as the target error or the level of overparameterization. The only two exceptions are by Ji and Telgarsky (2019) and Chen et al. (2020), allowing a constant stepsize smaller than $2/\beta$, where $\beta$ is the smoothness of the objective near initialization. In comparison, our NTK result (see Theorem 6) handles arbitrarily large stepsizes, in particular, it recovers the related result of Ji and Telgarsky (2019) when specializing the stepsize to a constant.
@@ 2 GD for Logistic Regression
Assumption 1 (Bounded and separable data).
@@
Theorem 1 (Bounds on risk and phase transition time).
@@
[…]
@@
The proof of Theorem 1 is deferred to Appendix A. Theorem 1 provides control over the average loss in the EoS phase, the last iterate loss in the stable phase, and the maximum length of the EoS phase. Note that GD might never enter the EoS phase. For instance, this happens when the stepsize is sufficiently small such that 2 holds for $\mathbf{w}_{s}=\mathbf{w}_{0}=0$. However, if GD enters the EoS phase at the beginning of the optimization, the algorithm must undergo a phase transition and switch to the stable phase after a finite time. We emphasize that Theorem 1 allows an arbitrarily large $\eta$.
Benefits of large stepsize.
Theorem 1 shows that a large stepsize improves optimization efficiency in two ways. First, when $\eta=\mathcal{O}(1)$, the EoS phase ends in finite time, so asymptotically GD stays in the stable phase and attains an $\tilde{\mathcal{O}}\big{(}1/(\eta t)\big{)}$ rate. In this case, a larger stepsize leads to a smaller constant factor. We also note that this rate can be sharp ignoring logarithmic factors according to the experimental results in Figures 1(b) and 1(d). Interestingly, recall the classical rate of GD for smooth convex optimization is $\mathcal{O}\big{(}1/(\eta t)\big{)}$ for $\eta<2/\beta$, where $\beta$ is the smoothness parameter (Nesterov, 2018, Theorem 2.1.14). Our results (nearly) match this rate but remove the $\eta<2/\beta$ condition for logistic regression, a special class of smooth convex optimization problems.
Second, Theorem 1 suggests that GD with a larger stepsize converges faster in the stable phase, but it takes more steps to exit the initial EoS phase. Picking a stepsize proportional to the total number of steps will balance these two effects and lead to an acceleration effect. This is formalized by the following corollary (the proof is deferred to Appendix B).
Corollary 2 (Acceleration of a large stepsize).
Under the same setup as in Theorem 1, consider GD with a given budget of $T$ steps, where $T\geq 120\max\{e,n\}/\gamma^{2}.$ Then for
@@
Corollary 2 shows that GD attains an $\tilde{\mathcal{O}}(1/T^{2})$ loss after $T$ steps with a large stepsize $\eta=\Theta(T)$. We point out that the final low loss is a consequence of the initial EoS phase. In particular, the next theorem (the proof is deferred to Appendix C) shows that in general, GD with $T$ steps of fixed stepsize must suffer an $\Omega(1/T)$ loss if it never enters the EoS phase.
Theorem 3 (Lower bound in the classical regime).
@@
We remark that the acceleration in Corollary 2 requires the budget to be at least linear in the sample size, that is, $T\geq\Omega(n)$. This requirement is rooted in the phase transition time bound in Theorem 1. We conjecture our phase transition time bound is improvable and the $T\geq\Omega(n)$ condition in Corollary 2 can be relaxed.
@@
We now showcase a powerful yet surprisingly simple technique for handling large stepsizes, which lies at the heart of the proof of Theorem 1. To begin with, consider a comparator $\mathbf{u}:=\mathbf{u}_{1}+\mathbf{u}_{2}$. Then from the definition of (GD), we have
@@
Let us choose $\mathbf{u}_{1}:=\mathbf{w}_{*}\cdot\ln(\gamma^{2}\eta t)/\gamma$, then we have $L(\mathbf{u}_{1})\leq 1/(\gamma^{2}\eta t)$ and $\|\mathbf{w}_{0}-\mathbf{u}\|=\|\mathbf{u}\|=\ln(\gamma^{2}\eta t)/\gamma+\eta/(2\gamma)$. Then (3) allows us to control the average loss and the norm of $\mathbf{w}_{t}$ under any stepsize $\eta>0$. We refer the reader to Appendix A for a complete proof of Theorem 1.
@@
When $\eta\geq\omega(1)$ compared to the total number of steps $T$ (for example, $\eta=\Theta(T)$ as in Corollary 2), existing implicit bias results (Soudry et al., 2018; Ji and Telgarsky, 2018) can no longer be applied. In particular, our bound on the norm of the GD iterates is as large as $\mathcal{O}(\eta)$ (see (3) or Lemma 8 in Appendix A), which prevents us from proving a good margin for the GD iterates. Nonetheless, as the final GD iterate attains low training error, the standard uniform-convergence theory is applicable to guarantee its generalization. We refer the reader to Proposition 15 in Appendix D for one such bound.

### v2

@@ Abstract
We consider gradient descent (GD) with a constant stepsize applied to logistic regression with linearly separable data, where the constant stepsize $\eta$ is so large that the loss initially oscillates. We show that GD exits this initial oscillatory phase rapidly — in $\mathcal{O}(\eta)$ steps, and subsequently achieves an $\tilde{\mathcal{O}}(1/(\eta t))$ convergence rate after $t$ additional steps. Our results imply that, given a budget of $T$ steps, GD can achieve an accelerated loss of $\tilde{\mathcal{O}}(1/T^{2})$ with an aggressive stepsize $\eta:=\Theta(T)$, without any use of momentum or variable stepsize schedulers. Our proof technique is versatile and also handles general classification loss functions (where exponential tails are needed for the $\tilde{\mathcal{O}}(1/T^{2})$ acceleration), nonlinear predictors in the neural tangent kernel regime, and online stochastic gradient descent (SGD) with a large stepsize, under suitable separability conditions.
keywords:
optimization; logistic regression; gradient descent; edge of stability; acceleration
@@ 1 Introduction
The classical theory for constant-stepsize GD assumes a small stepsize such that the induced loss decreases monotonically (Nesterov, 2018). For example, if $L(\mathbf{w})$ is $\beta$-smooth, then choosing $\eta<2/\beta$ guarantees $L(\mathbf{w}_{t})$ to decrease monotonically, a result known as the descent lemma (Nesterov, 2018, Section 1.2.3). Fruitful convergence results have been developed for small stepsize GD based on the descent lemma or its variants. However, little theory is known when GD is run with a large stepsize (for instance, $\eta>2/\beta$ for a $\beta$-smooth function), which induces non-monotonic loss values (exceptions will be discussed later in Section 1.1). Cohen et al. (2020) refer to this as the edge of stability (EoS) regime, and they further point out that GD usually operates in the EoS regime for trained neural networks to exhibit reasonable optimization or generalization performance (Wu and Ma, 2018; Cohen et al., 2020).
@@
Our first main result (Theorem 2.1) characterizes the dynamics of large stepsize GD for logistic regression with linearly separable data. We show that GD initially induces a non-monotonic loss (the EoS phase) and then exits this regime in finite time (phase transition), and afterwards the loss decreases monotonically (the stable phase). Specifically:
@@
    The Stable phase.  Finally, we show in the stable phase, the loss decreases monotonically at an $\tilde{\mathcal{O}}\big{(}1/(\eta t)\big{)}$ rate after $t$ steps. Figures 1 and 1 empirically verify that this rate is sharp asymptotically (as $t\to\infty$) ignoring logarithmic factors.
The above result immediately justifies the benefits of a large stepsize for improving optimization efficiency. First, a larger $\eta$ yields a smaller constant factor in the asymptotic bound, $\tilde{\mathcal{O}}(1/(\eta t))$. We highlight that $\eta$ can be an arbitrarily large constant, beyond the classical wisdom that constrains $\eta$ by the inverse smoothness (Nesterov, 2018). Second, given a budget of $T$ steps, GD attains an $\tilde{\mathcal{O}}(1/T^{2})$ loss when equipped with an aggressive stepsize $\eta:=\Theta(T)$ that optimally balances the steps spent in the EoS and stable phases (see Corollary 2.2). In contrast, we also provide an $\Omega(1/T)$ loss lower bound for constant-stepsize GD that does not enter the EoS phase (see Theorem 2.3). Our theory explains the empirical observations in Figure 1.
[…]
Figure 1: Behaviors of (GD) for logistic regression (1) with initialization $\mathbf{w}_{0}:=0$. Figures 1 and 1 are results on a toy dataset consisting of four samples, $\mathbf{x}_{1}=(1,0.2),\ y_{1}=1$, $\mathbf{x}_{2}=(-2,0.2),\ y_{2}=1$, $\mathbf{x}_{3}=(-1,-0.2),\ y_{3}=-1$, and $\mathbf{x}_{4}=(2,-0.2),\ y_{4}=-1$. Here, the stepsize $\eta=2^{2}$ is small, and $\eta\in\{2^{3},2^{4},2^{5}\}$ is large (depending on whether the loss decreases monotonically). Figures 1 and 1 are results on $1000$ samples from the MNIST dataset with labels “0” or “8”. Here, the stepsize $\eta=2^{-6}$ is small, and $\eta\in\{2^{-4},2^{-2},2^{0}\}$ is large. Figures 1 and 1 suggest that (GD) with a larger stepsize incurs a smaller loss when stabilized. Figures 1 and 1 suggest that $L(\mathbf{w}_{t})=\Theta(1/(\eta t))$ asymptotically as $t\to\infty$.
@@
  • General losses.  We relax the logistic loss in (1) to general classification loss functions under a set of conditions (see Assumption 3 and Theorem 4.2). Specifically, the EoS bound relies on the Lipschitzness of the loss, the stable phase bound relies on an additional self-boundedness condition, and the $\tilde{\mathcal{O}}(1/T^{2})$ acceleration relies on the loss having an exponential tail.
  • Nonlinear models.  We extend the linear predictor in (1) to a two-layer network in the neural tangent kernel (NTK, Jacot et al., 2018) regime (Theorem 4.2). To our knowledge, this is the first NTK result that accommodates an arbitrarily large stepsize (the minimum network width depends on the stepsize). For a constant stepsize, the minimum network width is polylogarithmic in the number of steps and the number of samples, recovering the main result of Ji and Telgarsky (2019) as a special case.
  • SGD.  Finally, we consider constant-stepsize online stochastic gradient descent (SGD) for logistic regression with separable data. We provide upper bounds on the population logistic loss and the population zero-one error (Theorem 3.1), allowing the stepsize to be $o(T)$ (where $T$ is the maximum number of steps) and arbitrarily large, respectively. These are in contrast to the typical SGD results in online smooth convex optimization that only allow small (even vanishing) stepsizes (see, for example, Hazan, 2022, Chapter 3).
Implication and limitations of our theory in practice.
The work by Cohen et al. (2020) empirically identifies two phenomena when training neural networks using GD with a large stepsize ($\eta$): progressive sharpening, where the sharpness increases until around $2/\eta$, and EoS, where the sharpness oscillates around $2/\eta$. Their definition of EoS and ours are equivalent for quadratic functions, and our definition highlights the unstable loss behaviors more explicitly for non-quadratic functions. Additionally, a transition from the EoS phase to the stable phase is observed in neural networks trained with cross-entropy loss (see Cohen et al., 2020, Appendix A). Our theory offers valuable insights into the EoS and stable phases but remains inconclusive regarding the progressive sharpening phenomenon.
@@ 1.1 Related Works
Empirically, training neural networks with GD (or large batch SGD) often requires a large stepsize to achieve reasonable optimization and generalization performance (see, for example, Hoffer et al., 2017; Goyal et al., 2017; Wu and Ma, 2018; Cohen et al., 2020), whereby the training loss exhibits non-monotonic behavior. Cohen et al. (2020) introduce the term edge of stability (EoS) to refer to this. The theory of EoS and large stepsize has been studied in various cases, such as one- or two-dimensional (nearly analytic) functions (Zhu et al., 2022; Ahn et al., 2023; Kreisler et al., 2023; Chen et al., 2023; Wang et al., 2023), scale-invariant networks (Lyu et al., 2022), diagonal linear networks (Even et al., 2023; Andriushchenko et al., 2023), and matrix factorization (Wang et al., 2022b; Chen and Bruna, 2023). Besides, there are some general theories for EoS (Kong and Tao, 2020; Ahn et al., 2022; Damian et al., 2022; Ma et al., 2022; Wang et al., 2022c; Lu et al., 2023), but they are built up on subtle assumptions. In comparison, we work in a natural classification setting under standard assumptions.
Among the literature of EoS, the work by Wu et al. (2023) is most relevant to us. They show the convergence of GD with an arbitrarily large stepsize for logistic regression with linearly separable and non-degenerate data. Our work significantly improves theirs. First, the constant factors in their bounds are exponential in the stepsize $\eta$. In comparison, we nail down the dependence on $\eta$ in each phase (see Theorem 2.1), establishing the $\tilde{\mathcal{O}}(1/(\eta t))$ asymptotic rate and the $\tilde{\mathcal{O}}(1/T^{2})$ acceleration with stepsize $\eta=\Theta(T)$ (see Corollary 2.2). So we explain the benefits of large stepsizes while they cannot. Second, they require the dataset to be non-degenerate such that the support vectors span the dataset (see their Assumption 3), while we do not need this condition. Finally, their results are specialized to GD and logistic regression, while our approach is versatile and also handles general classification losses, nonlinear predictors in the NTK regime, and SGD with a large stepsize.
@@
Malitsky and Mishchenko (2020) design an adaptive stepsize scheduler for GD for smooth convex optimization, where they show a convergence rate without an explicit dependence on the global smoothness parameter. Their analysis involves two consecutive GD steps, instead of using the classical descent lemma. Thus their analysis also allows non-monotonic loss. In comparison, we focus on GD with a large but constant stepsize, and our analysis relies on the separability condition of the dataset instead of variable stepsizes.
@@
The neural tangent kernel (NTK) is most prominently presented and named by Jacot et al. (2018), and similar concepts are used by subsequent works to analyze the early training dynamics of GD. In the majority of these works (see Du et al., 2018; Zou et al., 2018; Zou and Gu, 2019; Allen-Zhu et al., 2019, for example), the stepsize of GD is $o(1)$ and vanishes as a function of other parameters such as the target error or the level of overparameterization. The only two exceptions are by Ji and Telgarsky (2019) and Chen et al. (2020), allowing a constant stepsize smaller than $2/\beta$, where $\beta$ is the smoothness of the objective near initialization. In comparison, our NTK result (see Theorem 4.2) handles arbitrarily large stepsizes, in particular, it recovers the related result of Ji and Telgarsky (2019) when specializing the stepsize to a constant.
@@ 2 GD for Logistic Regression
Assumption 1 (Bounded and separable data)
@@
Theorem 2.1 (Bounds on risk and phase transition time).
@@
    $L(\mathbf{w}_{t})\leq\frac{2F(\mathbf{w}_{s})+\ln^{2}(\gamma^{2}\eta(t-s))}{\gamma^{2}\eta(t-s)},\quad\text{where}\ \ F(\mathbf{w}):=\frac{1}{n}\sum_{i=1}^{n}\exp(-y_{i}\mathbf{x}_{i}^{\top}\mathbf{w}).$
@@
The proof of Theorem 2.1 is deferred to Appendix B. Theorem 2.1 provides control over the average loss in the EoS phase, the last iterate loss in the stable phase, and the maximum length of the EoS phase. Note that GD might never enter the EoS phase. For instance, this happens when the stepsize is sufficiently small such that 2 holds for $\mathbf{w}_{s}=\mathbf{w}_{0}=0$. However, if GD enters the EoS phase at the beginning of the optimization, the algorithm must undergo a phase transition and switch to the stable phase after a finite time. We emphasize that Theorem 2.1 allows an arbitrarily large $\eta$.
Benefits of large stepsizes.
Theorem 2.1 shows that a large stepsize improves optimization efficiency in two ways. First, when $\eta=\mathcal{O}(1)$, the EoS phase ends in finite time, so asymptotically GD stays in the stable phase and attains an $\tilde{\mathcal{O}}\big{(}1/(\eta t)\big{)}$ rate. In this case, a larger stepsize leads to a smaller constant factor. We also note that this rate can be sharp ignoring logarithmic factors according to the experimental results in Figures 1 and 1. Interestingly, recall the classical rate of GD for smooth convex optimization is $\mathcal{O}\big{(}1/(\eta t)\big{)}$ for $\eta<2/\beta$, where $\beta$ is the smoothness parameter (Nesterov, 2018, Theorem 2.1.14). Our results (nearly) match this rate but remove the $\eta<2/\beta$ condition for logistic regression, a special class of smooth convex optimization problems.
Second, Theorem 2.1 suggests that GD with a larger stepsize converges faster in the stable phase, but it takes more steps to exit the initial EoS phase. Picking a stepsize proportional to the total number of steps will balance these two effects and lead to an acceleration effect. This is formalized by the following corollary (the proof is deferred to Appendix C).
Corollary 2.2 (Acceleration of a large stepsize).
Under the same setup as in Theorem 2.1, consider GD with a given budget of $T$ steps, where $T\geq 120\max\{e,n\}/\gamma^{2}.$ Then for
@@
Corollary 2.2 shows that GD attains an $\tilde{\mathcal{O}}(1/T^{2})$ loss after $T$ steps with a large stepsize $\eta=\Theta(T)$. We point out that the final low loss is a consequence of the initial EoS phase. In particular, the next theorem (the proof is deferred to Appendix D) shows that in general, GD with $T$ steps of fixed stepsize must suffer an $\Omega(1/T)$ loss if it never enters the EoS phase.
Theorem 2.3 (Lower bound in the classical regime).
@@
We remark that the acceleration in Corollary 2.2 requires the budget to be at least linear in the sample size, that is, $T\geq\Omega(n)$. This requirement is rooted in the phase transition time bound in Theorem 2.1. We conjecture our phase transition time bound, especially its dependence on the sample size, is improvable, and thus the $T\geq\Omega(n)$ condition in Corollary 2.2 can be relaxed. Some empirical evidence is provided in Figure 2 in Appendix A.
@@
We now showcase a powerful yet surprisingly simple technique for handling large stepsizes, which lies at the heart of the proof of Theorem 2.1. To begin with, consider a comparator $\mathbf{u}:=\mathbf{u}_{1}+\mathbf{u}_{2}$. Then from the definition of (GD), we have
@@
Let us choose $\mathbf{u}_{1}:=\mathbf{w}_{*}\cdot\ln(\gamma^{2}\eta t)/\gamma$, then we have $L(\mathbf{u}_{1})\leq 1/(\gamma^{2}\eta t)$ and $\|\mathbf{w}_{0}-\mathbf{u}\|=\|\mathbf{u}\|=\ln(\gamma^{2}\eta t)/\gamma+\eta/(2\gamma)$. Then (3) allows us to control the average loss and the norm of $\mathbf{w}_{t}$ under any stepsize $\eta>0$. We refer the reader to Appendix B for a complete proof of Theorem 2.1.
Other acceleration techniques.
Our focus in this paper is to understand the benefits of using a large but constant stepsize in GD. Based on a simple yet powerful technique, we show that GD with a suitably large stepsize leads to non-monotonic loss and achieves improved optimization efficiency. Our techniques are orthogonal to standard acceleration methods such as momentum and variable stepsize schedulers. We conjecture these techniques can be combined to obtain an even faster optimization. This is left for future work.
@@
When $\eta\geq\omega(1)$ compared to the total number of steps $T$ (for example, $\eta=\Theta(T)$ as in Corollary 2.2), existing implicit bias results (Soudry et al., 2018; Ji and Telgarsky, 2018) can no longer be applied. In particular, our bound on the norm of the GD iterates is as large as $\mathcal{O}(\eta)$ (see (3) or Lemma B.3 in Appendix B), which prevents us from proving a good margin for the GD iterates. Nonetheless, as the final GD iterate attains low training error, the standard uniform-convergence theory is applicable to guarantee its generalization. We refer the reader to Proposition E.1 in Appendix E for one such bound.